OpenStudy (anonymous):

One positive integer is 6 less than twice another. The sum of their squares is 377. Find the integer.

OpenStudy (anonymous):

start with \[x\] and \[2x-6\]

OpenStudy (anonymous):

then \[x^2+(2x-6)^2=377\] solve for x

OpenStudy (anonymous):

\[x^2+(2x-6)^2=377\] \[x^2+4x^2-24x+36=377\] \[5x^2-24x-341=0\] \[(x-11)(5x+31)=0\] etc

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (anonymous):

yw